# Burnside type results for fusion categories

### v10

@@ 2.3. Frobenius-Perron theory for a RN-hypergroups
Frobenius-Perron theory can be defined for RN-hypergroups, in the same manner as for fusion rings, see [EGNO15, §3]. We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$. As for fusion rings, by a similar argument, it can be shown that $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$. Indeed, it easy to see that the matrix $N_{i^{*}}$ corresponding to $x_{i^{*}}$ is the transpose of the matrix $N_{i}$ of $x_{i}$ multiplied by the scalars $\frac{h_{j}}{h_{k}}$ in the $(j,k)$-entry. More precisely, $(N_{i^{*}})_{jk}=\frac{h_{j}}{h_{k}}(N_{i})_{kj}$. Thus, the two matrices have the same set of eigenvalues, in particular the same Frobenius-Perron eigenvalue. Note that any RN-hypergroup ${(H,\mathcal{B})}$ can be normalized by $\mu_{1}=\mathrm{FPdim}$, since $\mu_{1}(x_{i})>0$ for all $1\leq i\leq m$.
@@
Note that any symmetric RN-hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1]. Indeed the function $\tau$ from §3.1 is symmetric and therefore $\tau(x_{i}x_{j}x_{k})=\tau(x_{j}x_{k}x_{i})=\tau(x_{k}x_{i}x_{j})$. Since $\tau(x_{i}x_{j}x_{k})=h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})$ one has
(2.16) $h_{i}^{-1}m(x_{i^{*}},x_{j}x_{k})=h_{j}^{-1}m(x_{j^{*}},x_{k}x_{i})=h_{k}^{-1}m(x_{k^{*}},x_{i}x_{j}).$
which shows that ${(H,\mathcal{B})}$ is transitive.
Then it is easy to see that an analogue [EGNO15, Proposition 3.3.6] also holds for symmetric RN-hypergroups. In particular it follows that $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique which takes positive values on $\mathcal{B}$.
@@
In this case of a RN-hypergroup ${(H,\mathcal{B})}$, one can define the Frobenius-Perron dimension as
$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
@@
For a normalized RN-hypergroup,
$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
@@
It is easy to see that [EGNO15, Proposition 3.3.6] also works for normalized RN-hypergroups instead of fusion rings. Therefore $\mu_{1}:H\rightarrow H,x_{i}\mapsto 1$ is the unique morphism $\phi:H\rightarrow\mathbb{C}$ with $\phi(x_{i})>0$ for all $1\leq i\leq m$. This implies that $\mathrm{FPdim}(x_{i})=1$ for all $1\leq i\leq m$ and therefore $n{(H,\mathcal{B},\mathrm{FPdim})}=\mathrm{FPdim}{(H,\mathcal{B})}$. ∎
Note that by Lemma 2.2, any RN-hypergroup is normalizable via the character $\mu_{1}=\mathrm{FPdim}$. Lemma 2.13 also shows that
$n(H,\mathcal{B},\mathrm{FPdim})=\sum_{i=1}^{m}h_{i}|\mu_{1}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}\mathrm{FPdim}(x_{i})^{2}=\mathrm{FPdim}{(H,\mathcal{B})}$
for any RN-hypergroup. For a RN-hypergroup, we sometimes write ${(H,\mathcal{B},\mathrm{FPdim})}$ instead of ${(H,\mathcal{B})}$ to specify that we normalize it by $\mathrm{FPdim}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be a normalizable hypergroup. Then the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling. If ${(H,\mathcal{B})}$ is RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is also invariant under rescaling with real positive numbers.
@@
First we show that the order of a normalizable hypergroup is invariant under rescaling. Let $y_{i}:=\frac{x_{i}}{\alpha_{{}_{i}}}$, for some non-zero complex numbers $\alpha_{{}_{i}}$ with $\alpha_{1}=1$, $\alpha_{{}_{i^{*}}}=\overline{\alpha_{{}_{i}}}$. Consider $(H,\mathcal{B}^{\prime},\mu_{1})$ the rescaled hypergroup with :=$\mathcal{B}^{\prime}:=\{y_{i}\}_{i\in\mathcal{I}}$. Note that the order of $y_{i}$ is given by ${h^{\prime}_{i}}^{-1}=N^{1^{\prime}}_{i^{\prime}i^{{}^{\prime}*}}=\tau(x_{i^{\prime}}x_{i^{{}^{\prime}*}})=\frac{1}{|\alpha_{{}_{i}}|^{2}}h_{i}^{-1}$ and this implies that $h^{\prime}_{i}=|\alpha_{{}_{i}}|^{2}h_{i}>0.$ Therefore
$n(H,\mathcal{B}^{\prime},\mu_{1})=\sum_{i=1}^{m}h^{\prime}_{i}\mid\mu_{1}(y_{i})\mid^{2}=\sum_{i=1}^{m}h_{i}\mid\mu_{1}(x_{i})\mid^{2}=n{(H,\mathcal{B},\mu_{1})}.$
Moreover, if ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $1\leq i\leq m$, it follows that
$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i=1}^{m}h^{\prime}_{i}\mathrm{FPdim}(y_{i})^{2}=$
$=$ $\sum_{i=1}^{m}\mid\alpha_{{}_{i}}\mid^{2}h_{i}(\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}})^{2}$
$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.$
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then with the above notations, the dual formal codegrees are given by
(2.17) $\widehat{n}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i}d_{i}d_{i^{*}}}.$
Proof.
[…]
Remark 2.16.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian rational normalizable hypergroup. In order to find the central primitive idempotents of $H^{*}$, one can apply Remark 2.1 after normalization. If $x_{i}^{\circ}\in H^{*}$ are defined as above by $\langle x_{i}^{\circ},x_{i^{\prime}}\rangle=\delta_{i,i^{\prime}}$, then ${\widehat{E}}_{i}:=d_{i}x_{i}^{\circ}\in H^{*}$ are the primitive central idempotents of $H^{*}$ corresponding to the character
$\begin{array}[]{cccc}\widetilde{\omega}_{i}=\operatorname{ev}_{{\frac{x_{i}}{d_{i}}}}:&H^{*}&\to&\mathbb{C}\\ &\mu_{j}&\mapsto&\mu_{j}({\frac{x_{i}}{d_{i}}}).\end{array}$
2.4. Grouplike elements in hypergroups
In the rest of this paper all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$ and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$.
Lemma 2.17.
For any symmetric RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
Proof.
One has
(2.18) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=1}^{m}N^{k}_{ii^{*}}x_{k}.$
Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=1}^{m}N^{k}_{ii^{*}}d_{k}=1$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$. Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}\mathrm{FPdim}(x_{i})^{2}=1$. ∎
Corollary 2.18.
In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=1$ if and only if $d_{i}=1$.
@@ 2.4. Grouplike elements in hypergroups
For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ one has that $x_{i}\in G{(H,\mathcal{B})}$ if and only if
(2.19) ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$
Indeed applying $\mu_{1}$ to Equation 1.1 one has $h_{i}d_{i}d_{i^{*}}=1$ and Equation (2.19) follows. The converse also follows by inspecting the multiplicity of $x_{1}$ in $x_{i}x_{i^{*}}$.
@@
(2.20) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{J}.$
@@
If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.19) one obtains $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$ since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$. Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$ then since ${(H,\mathcal{B})}$ is abelian it follows that ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=1$ and therefore $h_{i}x_{i}x_{i^{*}}=1$. ∎
Dually one has the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if
@@
The dual version of Equation (2.19) implies that $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Then evaluating at ${\frac{x_{i}}{d_{i}}}$ one gets the desired identity. ∎
@@
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ for any $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ one has that $n_{j}=n(H,B,\mu_{1})$.
@@
By Lemma 2.21 one has $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$. By Equation (2.5) one has $n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$. ∎
@@
(2.22) $|\mu_{j}(x_{i})|\leq|d_{i}|\;\;\text{for all}\;i\in\mathcal{I}\;\text{and}\;j\in\mathcal{J}.$
@@
Suppose first that ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$. Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$ in this case. One has $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$ and therefore $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$. It follows that $|\mu_{j}(x_{i})|\leq d_{i}$.
Suppose now that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.11 in this case $\mu_{j}\mapsto 1$ is the only morphism on ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$. Then, the equation
(2.23) $\mu_{j}=\sum_{i=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}$
@@
  1. (1)
    ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$
  2. (2)
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$ for all $j\in\mathcal{J}$. Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
@@
$n_{j}=\sum_{i=1}^{m}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i=1}^{m}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$
and equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$. By Lemma 2.21 this happens if and only if $\mu_{j}\in G{(H,\mathcal{B})}$. ∎

### v11

@@ 2.3. Frobenius-Perron theory for a RN-hypergroups
Frobenius-Perron theory can be defined for RN-hypergroups, in the same manner as for fusion rings, see [EGNO15, §3]. We denote by $\mathrm{FPdim}(x_{i})$ the Frobenius-Perron eigenvalue of the left multiplication operator by $x_{i}$ on $H$, i.e. the ($\ell^{2}$) matrix norm of $N_{i}=(N_{i,j}^{k})_{j,k}$. Recall that a hypergroup is called symmetric when $N_{a,b}^{1}=N_{b,a}^{1}$, for all $a,b\in\mathcal{B}$, but $N_{a,b}^{1}=\delta_{a^{*},b}h_{a}^{-1}$, so symmetric means that $h_{a}=h_{a^{*}}$, for all $a\in\mathcal{B}$
Lemma 2.14.
An abelian hypergroup is symmetric.
Proof.
By definition and abelian assumption, $h_{i}^{-1}:=N_{i,i^{*}}^{1}=N_{i^{*},i}^{1}=h_{i^{*}}^{-1}$. ∎
Lemma 2.15.
If $N_{i,j}^{k}$ is nonzero then
(2.16) $\frac{h_{i^{*}}h_{j^{*}}h_{k}}{h_{i}h_{j}h_{k^{*}}}=1.$
Proof.
The associativity of $H$ reformulates as $\sum_{s}N_{i,j}^{s}N_{s,k}^{t}=\sum_{s}N_{j,k}^{s}N_{i,s}^{t}$, for all $i,j,k,t$. If $t=1$, we get that $\sum_{s}N_{i,j}^{s}N_{s,k}^{1}=\sum_{s}N_{j,k}^{s}N_{i,s}^{1}$. But recall that $N_{a,b}^{1}=\delta_{a^{*},b}h_{a}^{-1}$. It follows that $N_{i,j}^{k^{*}}h_{k^{*}}^{-1}=N_{j,k}^{i^{*}}h_{i}^{-1}$, in other words, $N_{i,j}^{k}=\frac{h_{k}}{h_{i}}N_{j,k^{*}}^{i^{*}}$. By applying this last equality, we get that $N_{j,k^{*}}^{i^{*}}=\frac{h_{i^{*}}}{h_{j}}N_{k^{*},i}^{j^{*}}$, and $N_{k^{*},i}^{j^{*}}=\frac{h_{j^{*}}}{h_{k^{*}}}N_{i,j}^{k}$. Thus
(2.17) $N_{i,j}^{k}=\frac{h_{k}}{h_{i}}N_{j,k^{*}}^{i^{*}}=\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}N_{k^{*},i}^{j^{*}}=\frac{h_{i^{*}}h_{j^{*}}h_{k}}{h_{i}h_{j}h_{k^{*}}}N_{i,j}^{k}$
The result follows. ∎
Theorem 2.16.
A (finite-dimensional) hypergroup is symmetric, i.e. $h_{i^{*}}=h_{i}$, for all $i\in\mathcal{I}$. In particular, the function $\tau$ from §2.1 is symmetric.
Proof.
Introduce the notation $r_{i}:=\frac{h_{i^{*}}}{h_{i}}$. Lemma 2.15 reformulates as:
$N_{i,j}^{k}\neq 0\Rightarrow r_{i}r_{j}=r_{k}.$
We aim to demonstrate that $r_{i}$ is a root of unity for every $i\in\mathcal{I}$. First, select $k_{2}$ such that $N_{i,i}^{k_{2}}$ is nonzero, which implies $r_{i}^{2}=r_{k_{2}}$. Next, choose $k_{3}$ such that $N_{i,k_{2}}^{k_{3}}$ is nonzero, leading to $r_{i}r_{k_{2}}=r_{k_{3}}$ and thus $r_{i}^{3}=r_{k_{3}}$. Continuing this process, we find that the sequence $\{r_{i}^{n}\}$ for $n\in\mathbb{N}$ matches the set $\{r_{j}\}$ for $j\in\mathcal{J}$, with $\mathcal{J}$ being a subset of $\mathcal{I}$. Due to the finite dimensionality of the hypergroup, $\mathcal{I}$ is finite, which implies that $\mathcal{J}$ is also finite. Consequently, there exist integers $m$ and $n$ with $m>n$ such that $r_{i}^{m}=r_{i}^{n}$. Thus, $r_{i}^{s}=1$ where $s=m-n>0$, confirming that $r_{i}$ is a root of unity. Additionally, the axioms of the hypergroup state that $N_{a,a^{*}}^{1}$ is positive for all $a\in\mathcal{B}$; hence $h_{i}$ is positive for all $i\in\mathcal{I}$. Therefore, $r_{i}=\frac{h_{i^{*}}}{h_{i}}$ is a positive root of unity, which must be 111. ∎
In the rest of the paper, we assume that the involution on $\mathcal{B}$ extends into an anti-involution on $H$, i.e. $N_{i,j}^{k}=N_{j^{*},i^{*}}^{k^{*}}$, for all $i,j,k$.
Proposition 2.17 (Frobenius Reciprocity).
The following equalities hold:
(2.18) $h_{k}^{-1}N_{i,j}^{k}=h_{j}^{-1}N_{i^{*},k}^{j}=h_{i}^{-1}N_{j,k^{*}}^{i^{*}}=h_{k}^{-1}N_{j^{*},i^{*}}^{k^{*}}=h_{j}^{-1}N_{k^{*},i}^{j^{*}}=h_{i}^{-1}N_{k,j^{*}}^{i}.$
Proof.
By (2.17), $N_{i,j}^{k}=\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}N_{k^{*},i}^{j^{*}}$, then by anti-involution, $N_{k^{*},i}^{j^{*}}=N_{i^{*},k}^{j}$, and by (2.16), $\frac{h_{i^{*}}h_{k}}{h_{i}h_{j}}=\frac{h_{k^{*}}}{h_{j^{*}}}$. But $h_{k^{*}}=h_{k}$ and $h_{j^{*}}=h_{j}$ by Theorem 2.16. The first equality follows. The rest is similar. ∎
Corollary 2.18.
For every $x_{i}\in\mathcal{B}$ then $\mathrm{FPdim}(x_{i})=\mathrm{FPdim}(x_{i^{*}})$.
Proof.
Let $N_{i}$ be the matrix $(N_{i,j}^{k})_{j,k}$. By the first equality of (2.18), $N_{i,j}^{k}=\frac{h_{k}}{h_{j}}N_{i^{*},k}^{j}$, so $N_{i}=DN_{i^{*}}^{T}D^{-1}$, with $D={\rm diag}(h_{i})$ and $()^{T}$ the matrix transpose. Thus
$\mathrm{FPdim}(x_{i^{*}})=\|N_{i^{*}}\|=\|N_{i^{*}}^{T}\|=\|D^{-1}N_{i}D\|=\|N_{i}\|=\mathrm{FPdim}(x_{i}).$
The second last equality hold because $D^{-1}N_{i}D$ have the same eigenvalues than $N_{i}$, because if $N_{i}v=\lambda v$ then $D^{-1}N_{i}Dw=\lambda w$, with $w=D^{-1}v$. ∎
Since we are dealing with the finite-dimensional case, by Theorem 2.16, every hypergroup is symmetric. Therefore, we can omit this assumption in the rest of the paper, as it is automatically satisfied.
@@
Note that any hypergroup ${(H,\mathcal{B})}$ is transitive in the sense of [EGNO15, Definition 3.3.1], i.e. $\forall i,j\in\mathcal{I}$ there are $k_{1},k_{2}\in\mathcal{I}$ such that $N_{i,k_{1}}^{j}$ and $N_{k_{2},i}^{j}$ are nonzero. Indeed, take $k_{1},k_{2}$ such that $N_{i^{*},j}^{k_{1}}$ and $N_{j,i^{*}}^{k_{2}}$ are nonzero, the result follows by Frobenius reciprocity (2.18).
Observe that an analogue of [EGNO15, Proposition 3.3.6] holds for RN-hypergroups as it requires Frobenius-Perron [EGNO15, Theorem 3.2.1] involving a RN-matrix (i.e. with real nonnegative entries). In particular, $\mathrm{FPdim}:H\rightarrow\mathbb{C}$ is the unique algebra morphism which takes positive values on $\mathcal{B}$.
@@
The Frobenius-Perron dimension of a RN-hypergroup ${(H,\mathcal{B})}$ is
$\mathrm{FPdim}{(H,\mathcal{B})}:=\sum_{i\in\mathcal{I}}h_{i}\mathrm{FPdim}(x_{i})^{2}.$
By Lemma 2.3 and Remark 2.19, any RN-hypergroup is normalizable via the linear character $\mu_{1}=\mathrm{FPdim}$. By Definition 2.5 and positivity of $\mathrm{FPdim}$,
$\mathrm{FPdim}{(H,\mathcal{B})}=n{(H,\mathcal{B},\mathrm{FPdim})}.$
Moreover, in the normalized case, by uniqueness in Remark 2.19, $\mathrm{FPdim}$ must be the augmentation map, i.e. $\mathrm{FPdim}(x_{i})=1$ for all $i\in\mathcal{I}$.
Recall from Proposition 2.6 that for a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, the order $n{(H,\mathcal{B},\mu_{1})}$ is invariant under rescaling.
Proposition 2.21.
If ${(H,\mathcal{B})}$ is a RN-hypergroup, then the Frobenius-Perron dimension of ${(H,\mathcal{B})}$ is invariant under rescaling with real positive numbers.
Proof.
If ${(H,\mathcal{B})}$ is a RN-hypergroup and $\alpha_{{}_{i}}=\alpha_{{}_{i^{*}}}>0$, for all $i\in\mathcal{I}$, then $(H^{\prime},\mathcal{B}^{\prime})=(H,\{\frac{x_{i}}{\alpha_{i}}\})$ is RN and
$\mathrm{FPdim}(H^{\prime},\mathcal{B}^{\prime})$ $=$ $\sum_{i\in\mathcal{I}}h^{\prime}_{i}\mathrm{FPdim}(\frac{x_{i}}{\alpha_{i}})^{2}=$
$=$ $\sum_{i\in\mathcal{I}}|\alpha_{{}_{i}}|^{2}h_{i}\frac{\mathrm{FPdim}(x_{i})^{2}}{\alpha_{{}_{i}}^{2}}$
$=$ $\mathrm{FPdim}{(H,\mathcal{B})}.\qed$
2.4. Fourier transform
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Define
(2.19) $\mathcal{F}:{(H,\mathcal{B})}\rightarrow{\widehat{(H,\mathcal{B})}},\;x_{i}\mapsto\frac{n{(H,\mathcal{B},\mu_{1})}}{h_{i^{*}}}x_{i^{*}}^{\circ}.$
It is clear that $\mathcal{F}$ is a linear isomorphism. Additionally, as noted in Remark 2.11, we have $\widehat{F}_{i}:=d_{i}x_{i}^{\circ}$. Therefore:
(2.20) $\mathcal{F}(x_{i})=\frac{n{(H,\mathcal{B},\mu_{1})}}{d_{i^{*}}h_{i^{*}}}\widehat{F}_{i^{*}}.$
For the sake of brevity, we define $|H|:=n{(H,\mathcal{B},\mu_{1})}$.
@@ 2.4. Fourier transform
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. With the above notations, for all $x,y\in H$:
$\langle\mathcal{F}(y),x\rangle=|H|\tau(xy).$
@@
It suffices to verify the above identity for $(x,y)=(x_{i},x_{j})$, for all $i,j$. We have:
[…]
since $h_{i}=h_{i^{*}}$ (see Lemma 2.14), and $\delta_{i,j^{*}}h_{i}^{-1}=\tau(x_{i}x_{j})$, because $h_{i}^{-1}=N_{i,i^{*}}^{1}$ and $\tau=x_{1}^{\circ}$. ∎
@@
Using the above notations, we have that
(2.21) $\mathcal{F}(F_{k})=\frac{|H|}{n_{k}}\mu_{k}.$
@@
By taking $(x,y)=(F_{j},F_{k})$ in Lemma 2.22, we obtain:
$\langle\mathcal{F}(F_{k}),F_{j}\rangle=|H|\tau(F_{j}F_{k})=\delta_{j,k}|H|\tau(F_{j}).$
From Equation (2.4), we have:
$\tau(F_{j})=\frac{1}{n_{j}}\big{(}\sum_{i\in\mathcal{I}}h_{i}\mu_{j}(x_{i^{*}})\tau(x_{i})\big{)}=\frac{1}{n_{j}}.$
Thus
$\langle\mathcal{F}(F_{k}),F_{j}\rangle=\delta_{j,k}\frac{|H|}{n_{j}},$
which shows that $\mathcal{F}(F_{k})=\frac{|H|}{n_{k}}\mu_{k}$, since $\mu_{i}(F_{j})=\delta_{i,j}$. ∎
2.5. Grouplike elements in hypergroups
In the rest of this paper, all abelian RN-hypergroups are normalized by $\mu_{1}=\mathrm{FPdim}$, and therefore their duals are also considered with respect to $\mu_{1}=\mathrm{FPdim}$. Recall that $d_{i}:=\mu_{1}(x_{i})=\mathrm{FPdim}(x_{i})$.
Lemma 2.24.
For any RN-hypergroup ${(H,\mathcal{B})}$, $h_{i}d_{i}^{2}\geq 1$. Moreover, $h_{i}d_{i}^{2}=1$ if and only $h_{i}x_{i}x_{i^{*}}=x_{1}.$
Proof.
The following equality holds
(2.22) $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}+\sum_{k=2}^{m}N^{k}_{ii^{*}}x_{k}.$
Passing to $\mathrm{FPdim}$, $h_{i}d_{i}^{2}=1+h_{i}\big{(}\sum_{k=2}^{m}N^{k}_{ii^{*}}d_{k}\big{)}\geq 1.$ If $h_{i}d_{i}^{2}=1$ then $\sum_{k=2}^{m}N^{k}_{ii^{*}}d_{k}=0$, and therefore $N^{k}_{ii^{*}}=0$ for all $k\neq 1$. Thus $h_{i}x_{i}x_{i^{*}}=x_{1}.$ Conversely, if $h_{i}x_{i}x_{i^{*}}=x_{1}$, applying $\mathrm{FPdim}$, it follows that $h_{i}d_{i}^{2}=1$. ∎
@@ 2.5. Grouplike elements in hypergroups
In any fusion ring ${(H,\mathcal{B})}$, $x_{i}x_{i^{*}}=x_{1}$ if and only if $d_{i}=1$.
@@
For a normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, we establish that $x_{i}\in G{(H,\mathcal{B})}$ if and only if:
(2.23) $\frac{x_{i}x_{i^{*}}}{d_{i}d_{i^{*}}}=x_{1}.$
Applying $\mu_{1}$ to Equation (1.1), we find that $h_{i}d_{i}d_{i^{*}}=1$, from which Equation (2.23) follows. The converse is similar.
@@
(2.24) $|\mu_{j}({\frac{x_{i}}{d_{i}}})|=1\;\;\text{for all}\;\;j\in\mathcal{I}.$
@@
If $x_{i}\in G{(H,\mathcal{B})}$ then applying $\mu_{j}$ to Equation (2.23) results in $|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1$, since $\mu_{j}(x_{i^{*}})=\overline{\mu_{j}(x_{i})}$. Conversely, if $\mu_{j}({\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}})=1$ for all $\mu_{j}$, then ${\frac{x_{i}}{d_{i}}}\frac{x_{i^{*}}}{d_{i^{*}}}=x_{1}$, since ${(H,\mathcal{B})}$ is abelian, and therefore $x_{i}\in G{(H,\mathcal{B})}$ by Remark 2.26. ∎
Dually we have the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ if and only if
@@
The dual version of Equation (2.23) implies that $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ if and only if $\mu_{j}\mu_{j^{\#}}=\mu_{1}.$ Evaluating at ${\frac{x_{i}}{d_{i}}}$ results in the desired identity. ∎
@@
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, for any $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, it holds that $n_{j}=n(H,B,\mu_{1})$.
@@
By Lemma 2.28, we find that $\mid\mu_{j}(x_{i})|=|d_{i}|$ for all $i\in\mathcal{I}$. From Equation (2.5), we deduce that $n_{j}=\sum_{i\in\mathcal{I}}h_{i}|\mu_{j}(x_{i})|^{2}=\sum_{i\in\mathcal{I}}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1})$. ∎
@@
(2.26) $|\mu_{j}(x_{i})|\leq|\mu_{1}(x_{i})|=|d_{i}|\;\;\text{, for all}\;i,j\in\mathcal{I}.$
@@
Suppose first that ${(H,\mathcal{B})}$ is RN and $\mu_{1}=\mathrm{FPdim}$. Thus $d_{i}={\mathrm{FPdim}(x_{i})}>0$. But $x_{i}=\sum_{j\in\mathcal{I}}\mu_{j}(x_{i})F_{j}$, and therefore $x_{i}F_{j}=\mu_{j}(x_{i})F_{j}$ meaning that $\mu_{j}(x_{i})$ is an eigenvalue of $L_{x_{i}}$. It follows that $|\mu_{j}(x_{i})|\leq d_{i}$, by definition of $\mathrm{FPdim}$.
Now, suppose that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. It follows that $\mathrm{FPdim}(\mu_{j})=1$, since by Remark 2.19, in this case, $\mu_{j}\mapsto 1$ is the only algebra morphism on $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ which takes positive values on $^{\widehat{\mathcal{B}}}$. Then, the equation
(2.27) $\mu_{j}=\sum_{i\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}$
@@
Under the same assumptions as Lemma 2.30, we have $n_{j}\leq n{(H,\mathcal{B},\mu_{1})}$, for all $j\in\mathcal{I}$. Moreover, $n_{j}=n{(H,\mathcal{B},\mu_{1})}$ if and only if $\mu_{j}$ is a grouplike element.
@@
$n_{j}=\sum_{i\in\mathcal{I}}h_{i}|\mu_{j}(x_{i})|^{2}\leq\sum_{i\in\mathcal{I}}h_{i}|d_{i}|^{2}=n(H,B,\mu_{1}),$
and the equality holds if and only if $|\mu_{j}(x_{i})|^{2}=|d_{i}|^{2}$, if and only if $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, by Lemma 2.28. ∎